OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 2x - cos 2x

OpenStudy (anonymous):

sin2x=2sinx*cox cos2x=cosx*cox-sinx*sinx also use identity...: 2 cos^2(x) - 1 = 1 - 2 sin^2(x) you should be abe to get it

OpenStudy (anonymous):

I don'y know how to answer it

OpenStudy (anonymous):

would you like me to give you the answer?

OpenStudy (anonymous):

@Rezz5 yes